# Hao Chen1, Bo Zhang1, Chunhua Shen1 1

### v1

@@ 3.5 Keyframe-based Video Generation
After obtaining the keyframes in the movie, we can generate video clips of the movie based on these keyframes. A straightforward approach is to utilize an existing image-to-video model, such as Stable Video Diffusion (SVD) [2], to generate these clips. Specifically, SVD transforms the input image into latent features for conditioning and introduces interaction with the CLIP features of the input image via cross-attention. Although SVD is capable of generating high-quality short videos, e.g., 25 frames, it struggles to generate longer movie clips.
@@
Figure 3: The ID-preserving renderer significantly enhances the perceived identity of the character.
@@ 4.1 Implementation Details
We use the U-Net of pretrained SDXL [24] as our diffusion decoder. The encoder consists of a pretrained CLIP image encoder (ViT-L) and a transformer-based compressor. The autoregressive model is initialized with a pretrained LLaMA-7B [36] checkpoint. All MLPs used to align different modal inputs follow the architecture in LLaVA [20]. Please refer to Appendix A for more details.
@@ 4.2 Experimental Setup
To systematically evaluate the effectiveness of our method and conduct comparisons, we construct a test dataset consisting of 10 long movies that are NOT included in the training set, with 50,000 keyframes after pre-processing. Each of these videos is processed through our proposed data pre-processing pipeline, which accurately extracts and annotates keyframes. For each keyframe, we generate corresponding textual prompts and character face embeddings.
@@
We evaluate the effectiveness of various methods using the following metrics. The CLIP score assesses semantic alignment between the plot and generated outputs. Quality is evaluated using the Aesthetic Score (AS), Frechet Image Distance (FID), and Inception Score (IS). For assessing consistency, we introduce metrics for short-term (ST) and long-term (LT) consistency. ST is measured by calculating the CLIP cosine similarity between adjacent frames containing the target character. LT consistency is evaluated by checking the presence of the target character in selected frames across the test set. Finally, we calculate the ratio of frames containing the target character to the total number of selected frames as a measure of the accuracy in maintaining long-term character consistency.
@@
Figure 4: Qualitative comparisons of story generation. Please refer to appendix for more comparisons.
@@ 4.3 Comparison with the State-of-the-Art
Many existing story generation methods focus on fine-tuning with small datasets, exhibiting poor generalization. Consequently, we compare our method only with those that demonstrate high generalization capabilities, namely StoryDiffusion [49] and StoryGen [19]. As shown in Figure 4, StoryDiffusion fails to maintain long-term consistency. For example, the character’s hair is inconsistent and the character in keyframe 77 is not consistent with keyframe 968. Similarly, StoryGen fails to preserve consistency and generates abnormal results. By contrast, our method achieves generating extremely long content while preserving both short-term and long-term consistency across multiple characters. The observation is also evidenced by the quantitative results in Table 1, where our method achieves high scores in both LT and ST. Moreover, the higher CLIP score reflects that our generated results align well with the storyline. Better IS, AS, and FID scores demonstrate our method achieves high-quality images. We showcase more qualitative results in Appendix D.
@@
Figure 5: Ablations on face embedding. Consistency is well-preserved with face embedding.
Table 1: Quantitative comparisons of different methods. ST and LT refer to our proposed short-term and long-term consistency metrics, respectively.
@@
StoryDiffusion 14.411 8.515 6.106 4.643 0.597 0.612
StoryGen 13.378 6.432 3.791 8.644 0.514 0.551
Ours 19.112 9.709 6.042 2.012 0.655 0.836
Ours-ref 19.972 9.800 6.096 1.879 0.669 0.851
@@
We perform a detailed comparison between our method and existing methods for generating long videos. For text-to-video methods, we use detailed descriptions prepared in the test set as input. For image-to-video approaches, we employ keyframes generated by our methods as inputs. As illustrated in Table 2, our method significantly outperforms prior open-source models in terms of quality, demonstrating strong generalization capabilities. Most importantly, our method is capable of generating videos that last hours with little compromise in quality, achieving state-of-the-art results. More qualitative video results are presented in the Appendix.
@@ 4.4 Analysis
Figure 6: After resolving the overfitting, our model is capable of generating more diverse results.
@@
The multi-modal script introduces face embedding to better preserve the consistency. Figure 5 compellingly demonstrates the effectiveness of this design. Specifically, the removal of face embeddings leads to a decreased ability of the model to preserve character consistency. Face embeddings carry more nuanced and precise information compared to text alone. With face embedding, both the short-term and long-term consistency are well-preserved.
@@ 5 Conclusion
We present MovieDreamer to address the challenge of generating long-duration visual content with complex narratives. This method elegantly marries the advantages of autoregression and diffusion, and is capable of generating long videos. Additionally, we design the multi-modal script which aims at preserving character consistency across generated sequences. We further introduce ID-preserving rendering to better preserve character IDs and support few-shot movie creation due to the in-context modeling. This work potentially opens up exciting possibilities for future advancements in automated long-duration video production.
@@
Figure 7: Our model is capable of generating results with better ID in few-shot scenario.
@@ Appendix A Implementation Details
We use the pretrained CLIP image encoder (ViT-S) to generate spatial tokens. The spatial tokens are compressed into two tokens using a network with 6 Transformer blocks. The architecture of compressor is shown in Figure 8. First, we process the spatial tokens using three transformers. We then interchange the dimensions of the hidden states and compress them using a linear layer. The compressed hidden states are further processed through three transformers to obtain the final compressed token. We find that setting the compressed token to 2 is sufficient for accurately reconstructing the image.
Subsequently, we employ the pretrained U-Net of SDXL[24] as our decoder to render the compressed tokens back into images. Notably, the original SDXL architecture demands an additional projected text embedding as input. To address this, we concatenate and project the compressed token into the desired input space of SDXL simply using a linear layer.
Compared to the image dataset, our collected movie dataset may lack diversity in objects, which somewhat limits the model’s performance. To ensure the model develops a general understanding of various objects first, we employed a two-stage training strategy to train a more effective model. In the first stage, we fine-tuned the pre-trained U-Net on a large existing image dataset at a resolution of 768 x 768, allowing the model to understand various objects. Our dataset comprises images from OpenImages[15], JourneyDB[32], and Object365[30]. In the second stage, we further fine-tuned the model using our self-collected movie dataset at a resolution of 896 x 512. The model is trained using AdamW, with a learning rate of 1e-4 scheduled by a cosine scheduler for 120k steps in the first stage, and a constant learning rate of 2e-5 for 120k steps in the second stage. For each stage, we first freeze the decoder and only train the compressor for 20k steps, driving it to compress images more appropriately. To enable the model to generate results consistent with movie lighting, we further set the noise offset to 0.05 during training. The entire training process of compressor and decoder takes 6 NVIDIA A800 GPUs.
@@
We utilize a pretrained LLaMA 7B model as the backbone for our autoregressive model, which is subsequently trained with our curated multi-modal script data. The data consists of 50,000 long videos with corresponding script annotations, which subsequently results in 5 million keyframes. We use two-layer MLPs with GELU as the activation function to unify different modalities into the input space of the large language model. We set the length of the context frames to 128. The global autoregressive model is trained for 3 days using 4 NVIDIA H100 GPUs with a constant learning rate of 2e-5 and 2k steps for warm-up.
To help the model distinguish the script for different frames, we inserted two special tokens, <START_OF_FRAME> and <END_OF_FRAME>, at the beginning and end of each frame’s script, respectively. Furthermore, in our setting, there are actually three different modalities: character ID, target frame, and detailed description. The character ID embedding is extracted with FARL [48], and for each frame we extract up to 5 characters. We encode target frame into compressed tokens with our trained encoder $\mathcal{E}$. The detailed description is divided into different sentences, and each sentence is encoded into one [CLS] token using LongCLIP [45]. We treat the remaining text such as "Characters:" and "Scene Elements:" as identifiers, which are processed same as LLaMA, as shown in Figure 2.
@@
Figure 13: Additional story comparison results. Our method is the first to generate extremely long contents while maintaining both short-term and long-term consistency across multiple characters. The model did not utilize any inputs from the training set during testing.
@@
To enhance our model’s ability to maintain character consistency, we additionally input the text embeddings encoded by LongCLIP and the character IDs extracted by FARL into the trained decoder. Both the text embeddings and character IDs are aligned with the decoder’s input using two-layer MLPs, and then concatenated with the compressed tokens to serve as the input for cross-attention.
We assume a maximum of 15 sentences for text embeddings and 5 character IDs. If the inputs do not meet these quantities, we pad them with zeros. Moreover, for the concatenated embeddings, we randomly mask each token with a probability of 0.2. This encourages the model to learn the relationship between facial features, text descriptions, and the target image.
@@
The $k$-mixutre GMM models the distribution $p(\mathbf{e}_{t}|\mathbf{H}_{<t},\mathbf{M}_{t})$, where $\mathbf{e}_{t}$ is the continuous real valued image tokens of sequence length $S$ and dimension $d$. $\mathbf{H}_{<t}$ and $\mathbf{M}_{t}$ refer to the previous history information and the current multi-modal script. The GMM consists of $k$ multivariate Gaussian distribution components that are independent across the channel dimension $d$. To construct the GMM, we replace the logit prediction layer of our autoregressive model initialized using LLaMA with a linear layer with $2kd+k$ dimension for output. Subsequently, the mean $m$ and the variance $v$ of shape $S\times kd$ and mixture weights $w$ of shape $S\times k$ are derived. $m$ consists of a stack of tensors $[m^{(1)},m^{(2)},\dots m^{(k)}]$, where $m_{i}$ is of shape $S\times d$. Similarly, $v=[v^{(1)},v^{(2)},\dots v^{(k)}]$. The mixture weight is normalized along the channel dimension. Given the ground truth image embedding $\mathbf{e}$, we aim to minimize its negative log likelihood within the GMM:
@@
During training, we additionally introduce the $\ell_{1}$ and $\ell_{2}$ loss between the predicted image tokens and the ground truth tokens to improve model’s performance. However, we find that these losses are not on the same scale as the negative log-likelihood. The negative log-likelihood is way larger than the $\ell_{1}$ and $\ell_{2}$ loss, which reduces the influence of the $\ell_{1}$ and $\ell_{2}$ loss during the optimization. The difference in scale is because the constant term and covariance matrix of the negative log-likelihood are directly related to the dimension $d$. Concretly, for a variable $\mathbf{x}$ with mean $\mu$, the calculation of the negative log-likelihood is essentially represented as:
@@
where $\frac{d}{2}\log(2\pi)$ is the constant term and $\Sigma$ is the covariance matrix. As $d$ increases, both the constant term and the size of the covariance matrix grow, resulting in significant increase of the results. To bring the different losses to the same scale, we scale the negative log-likelihood loss by dividing it by the dimension $d$. Consequently, the negative log-likelihood loss in Section 3.3 is formulated as:

### v2

@@ 3.5 Keyframe-based Video Generation
After obtaining the keyframes, we can generate video clips of the movie based on these keyframes. A straightforward approach is to utilize an existing image-to-video model, such as Stable Video Diffusion (SVD) (Blattmann et al., 2023), to generate these clips. Specifically, SVD transforms the input image into latent features for conditioning and introduces interaction with the CLIP features of the input image via cross-attention. Although SVD is capable of generating high-quality short videos, e.g., 25 frames, it struggles to generate longer movie clips.
@@
We want to emphasize that our key contribution lies in the hierarchical controllable very long video generation, with MLLM for keyframe generation and diffusion model for video rendering. Our subsequent experiments also demonstrate that existing high-quality video generation models trained by resource-rich institutions can be integrated with our MovieDreamer, producing better visually appealing results. This further demonstrates our flexibility.
@@
Figure 3: The ID-preserving renderer significantly enhances the perceived identity of the character. For the second row, the input is only the compressed token. For the third row, the input is the compressed token, face ID, and description embedding.
@@ 4.1 Implementation Details
We use pretrained models to reduce the training computational overhead, which is very common in the era of large models. We use the U-Net of pretrained SDXL (Podell et al., 2023) as our diffusion decoder. The encoder consists of a pretrained CLIP image encoder (ViT-L) (Radford et al., 2021) and a transformer-based compressor. The autoregressive model is initialized with a pretrained LLaMA-7B (Touvron et al., 2023) checkpoint. All MLPs used to align different modal inputs follow the architecture in LLaVA (Liu et al., 2023b). Please refer to Appendix A for more details about the network architecture and training.
@@ 4.2 Experimental Setup
The majority of our training data is sourced from Condensed Movies (Bain et al., 2020), with additional data collected using the same methodology as this dataset, resulting in 5M keyframes with corresponding script annotations. To systematically evaluate the effectiveness of our method, we construct a test dataset consisting of 100 long movies that are NOT included in the training set, with 1M keyframes after pre-processing. Each of these videos is processed through our proposed data pre-processing pipeline (see Appendix C), which accurately extracts and annotates keyframes.
@@
We evaluate the effectiveness of various methods using the following metrics. The CLIP score (Radford et al., 2021) assesses semantic alignment between the plot and generated outputs. Quality is evaluated using the Aesthetic Score (AS) (Schuhmann et al., 2022), Fréchet Image Distance (FID) (Heusel et al., 2018), and Inception Score (IS) (Salimans et al., 2016). For assessing consistency, we propose new metrics for short-term (ST) and long-term (LT) consistency. ST is measured by calculating the CLIP cosine similarity between adjacent frames containing the target character. LT consistency is evaluated by checking the presence of the target character in selected frames across the test set. Finally, we calculate the ratio of frames containing the target character to the total number of selected frames as a measure of the accuracy in maintaining long-term character consistency. Please refer to the Appendix B for more details about the evaluation metrics.
@@
Figure 4: Qualitative comparisons of story generation. Please refer to the appendix E for more comparisons. Each keyframe is generated with the corresponding multimodal script as input.
@@ 4.3 Comparison with the State-of-the-Art
Many existing story generation methods focus on fine-tuning with small datasets, exhibiting poor generalization. Consequently, we compare our method only with those that demonstrate high generalization capabilities, namely StoryDiffusion (Zhou et al., 2024) and StoryGen (Liu et al., 2023a). As shown in Figure 4, StoryDiffusion consumes significant computational overhead and tends to generate generic faces that do not align with the desired character. Similarly, StoryGen fails to preserve consistency and generates abnormal results. In contrast, our method achieves generating extremely long content while preserving both short-term and long-term consistency across multiple characters. The observation is also evidenced by the quantitative results in Table 1, where our method achieves high scores in both LT and ST. Moreover, the higher CLIP score reflects that our generated results align well with the storyline. Better IS, AS, and FID scores demonstrate our method achieves high-quality images. We showcase more qualitative results in Appendix E.
@@
Figure 5: Ablations on face embedding. Character consistency is well-preserved with face embedding. The autoregressive model generates compressed tokens with multimodal script as input. The results are rendered by the decoder $\mathcal{D}$, with compressed tokens as well as face embedding and description embedding as input.
Table 1: Quantitative comparisons of different methods. ST and LT refer to our proposed short-term and long-term consistency metrics, respectively. noC denotes without continuous token supervision.
@@
StoryDiffusion 15.232 8.739 6.134 4.643 0.627 0.596
StoryGen 13.261 6.216 3.837 8.557 0.508 0.542
Ours 19.584 9.698 6.093 2.043 0.646 0.814
Ours-ref 20.071 9.842 6.288 1.912 0.701 0.893
Ours-noC 18.833 8.609 5.893 2.714 0.606 0.755
Ours-noC-ref 19.431 8.774 5.979 2.560 0.616 0.776
@@
We perform a detailed comparison between our method and existing methods for generating long videos. For text-to-video methods, we use detailed descriptions prepared in the test set as input. For image-to-video approaches, we employ keyframes generated by our methods as inputs. As illustrated in Table 2 and Figure 6 (b), our method significantly outperforms prior open-source models in terms of quality, demonstrating strong generalization capabilities. Most importantly, our method is capable of generating videos that last tens of minutes or even hours with little compromise in quality, achieving state-of-the-art results. Here, we compare existing methods with our enhanced video model. As shown in Appendix D, when integrated with enterprise-level closed-source video models, our approach demonstrates even better results.
@@ 4.4 Analysis
Continuous token supervision.
We noticed a significant decline in the model’s ability to maintain ID consistency on the test set when using only L1 and L2 supervision. We believe this is because L1 and L2 losses are not good at modeling probability distributions, as they are more suitable for regression tasks. However, we need to model the complex conditional distribution of the autoregressive process. Therefore, we applied more refined supervision on continuous real-valued tokens. As illustrated in Figure 6, with continuous token supervision, the model is better capable of preserving target ID. Both the quality of the generated image and the preservation of character ID degenerate without continuous token supervision.
Figure 6: (a) With continuous token supervision, our model demonstrates stronger generalization capabilities and achieves higher-quality results. Better character ID is preserved with continuous token supervision. (b) Visualization of long video generated by different methods.
@@
Figure 7: After resolving the overfitting, our model is capable of generating more diverse results. Each keyframe is generated with the corresponding multimodal script as input.
@@
The multi-modal script introduces face embedding to better preserve consistency. Figure 5 compellingly demonstrates the effectiveness of this design. Specifically, the removal of face embeddings leads to a significantly decreased ability of the model to preserve character consistency. Face embeddings carry more nuanced and precise information compared to text alone. With face embedding, both the short-term and long-term consistency are well-preserved.
@@ 5 Conclusion
We present MovieDreamer to address the challenge of generating long-duration visual content with complex narratives. This method elegantly marries the advantages of autoregression and diffusion, and is capable of generating very long videos. Additionally, we design the multi-modal script which aims at preserving character consistency across generated sequences. We further introduce ID-preserving rendering to better preserve character IDs and support few-shot movie creation due to the in-context modeling. This work potentially opens up exciting possibilities for future advancements in automated long-duration video production.
@@
Figure 8: Our zero-shot generation already achieves high success rates and quality. However, it occasionally struggles to preserve ID consistency accurately. Few-shot generation can produce results with better ID preservation.
@@ Appendix A Implementation Details
We use the pretrained CLIP image encoder (Radford et al., 2021) to generate spatial tokens. The spatial tokens are compressed into two tokens using the compressor, a network with 6 Transformer blocks. The architecture of the compressor is shown in Figure 9. First, we process the spatial tokens using three transformer blocks. We then interchange the dimensions of the hidden states and compress them using a linear layer. The compressed hidden states are further processed through three transformers to obtain the final compressed token. We find that setting the compressed token to 2 is sufficient for well reconstructing the image.
Subsequently, we employ the pretrained U-Net of SDXL (Podell et al., 2023) as our decoder to render the compressed tokens back into images. Notably, the original SDXL architecture demands an additional projected text embedding as input. To address this, we concatenate and project the compressed token into the desired input space of SDXL simply using a linear layer.
Compared to the image dataset, our collected movie dataset may lack diversity in all kinds of objects in the world, which somewhat limits the model’s performance. To ensure the model develops a general understanding of various objects first, we employed a three-stage training strategy to train a more effective model. In the first stage, we fine-tuned the pre-trained U-Net and the compressor on the large existing image dataset at a resolution of 768 x 768 with a cosine scheduling learning rate that ranges from 1e-4 to 2e-5 and an AdamW optimizer for 120k steps, allowing the model to understand various objects. Our dataset comprises images from OpenImages (Kuznetsova et al., 2018), JourneyDB (Sun et al., 2023a), and Object365 (Shao et al., 2019).
In the second stage, we further fine-tuned the model using our self-collected movie dataset at a resolution of 896 x 512. The model is trained using AdamW with a constant learning rate of 2e-5 for 120k steps in the second stage. For the first stage and second stage, we first freeze the decoder and only train the compressor for the first 20k steps, driving it to compress images more appropriately. To enable the model to generate results consistent with movie lighting, we further set the noise offset to 0.05 during training.
In the third stage, to better preserve character ID, we freeze the compressor and concatenate the compressed tokens with the text embedding and face embedding as inputs to the cross-attention and further finetune the decoder. Text embedding and face embedding are available in multimodal script data. Since the dimensions of the description embedding and facial embedding differ from those of the compressed token, we first project them to the same dimension as the compressed token using 2 two-layer MLPs with GELU as the activation function. Given that different frames have varying numbers of descriptions and face IDs, we pad the projected description embedding and projected facial embedding with zero tensors to extend the sequence length to 15 and 5, respectively. For instance, if a frame has a projected description embedding of shape [12, D] and a projected facial embedding of shape [3, D], we pad them with zero tensors of shape [3, D] and [2, D] to achieve shapes of [15, D] and [5, D]. Finally, we concatenate the compressed tokens with the projected description embedding and projected facial embedding, resulting in a [22, D] tensor, which serves as the key and value input for the UNet cross-attention block. The decoder and the MLPs are trained using AdamW with a constant learning rate of 2e-5 for 120k steps and a noise offset of 0.05. The entire training process of the compressor and decoder takes 3 weeks with 6 NVIDIA A800 GPUs.
@@
We utilize a pretrained LLaMA 7B model as the backbone for our autoregressive model, which is subsequently trained with our curated multi-modal script data. We use two-layer MLPs with GELU as the activation function to unify different modalities into the input space of the large language model. We have three distinct modalities that need to be unified into the input space: compressed tokens, facial embeddings, and description embeddings. Therefore, three MLPs are trained alongside the autoregressive model. We set the length of the context frames to 128, which results in the max sequence length around 5000. The global autoregressive model is trained for 3 days using 4 NVIDIA H100 GPUs with a constant learning rate of 2e-5 and 2k steps for warm-up.
To help the model distinguish the script for different frames, we inserted two special tokens, [START_OF_FRAME] and [END_OF_FRAME], at the beginning and end of each frame’s script, respectively. Furthermore, in our setting, there are three different modalities: face embedding, compressed tokens, and description embedding. The face embedding is extracted with FARL (Zheng et al., 2022), and for each frame, we extract the top 5 characters. The compressed tokens are generated by leveraging our trained encoder $\mathcal{E}$, with the target frame as input. The description is divided into different sentences, and each sentence is encoded into one [CLS] token using LongCLIP (Zhang et al., 2024). We treat the remaining text such as ”Characters:” and ”Scene Elements:” as identifiers, which are tokenized the same as LLaMA, as shown in Figure 2.
The identifier and descriptive texts in multimodal script.
Here, we provide an example to help readers understand the multimodal script in detail. Consider the example multimodal script in Figure 2, [Characters:], [Scene Elements: ], [Plots:] and the symbol [-] are identifiers, while [A man in a black suit…], [A woman with a teal headdress…], [Crowded event], [Elegent Lighting], [A man … a woman … are dancing closely…] are text descriptions. As shown in Figure 2, orange parts are identifiers, and blue parts are text descriptions. The identifier structures the input format, aiding the autoregressive model in understanding the different types of content within the script. The description is responsible for providing detailed information about the Character, Scene Elements, and Plot.
@@
  In a distant northern land, there was a tribe renowned for its berserkers, known for their courage and fearlessness. The berserkers spread war across many lands. Aron was always at the forefront; he was both the chief and the fiercest warrior.
@@
However, one day, Aron discovered that their former victims had united to seek vengeance against their tribe. The war inflicted immense losses on their tribe. The tribe members all became aware of the sins they had committed in the past.
@@
Nevertheless, Aron clung to his previous war strategies. His wife, Eleanor was injured seriously during the escape. She urged him to see that war was merciless, with no victors—only suffering. Knowing Aron’s stubborn nature, Eleanor lay down in disappointment.
@@
Eleanor’s injury and the prolonged suffering eventually made Aron realize that Eleanor was right. He struggled to sleep each night. One morning, he stared into the distance, deep in thought. After much reflection, he decided to leave the tribe and embark on a journey of atonement alone.
@@
Gatsby often held incredibly lavish parties at his home. Many famous big shots would attend the parties, but people didn’t know who Gatsby really was. However, Nick received an invitation to Gatsby’s party.
@@
At the party, he met Gatsby, who wanted to rekindle his lost love with Nick’s cousin, Daisy. Nick agreed to help Gatsby, so Gatsby had many people meticulously decorate Nick’s house. As the appointed time approached, Gatsby became very nervous.
@@
Daisy arrived as promised. Seeing Daisy, Gatsby’s long-suppressed feelings burst forth. Daisy was also thrilled to see Gatsby. They reignited their past romance.
@@
Gatsby made Daisy and her husband, Tom, into a room for a confrontation, hoping Daisy would leave with him. However, Tom had already investigated Gatsby’s background. He began to publicly expose Gatsby. At first, Gatsby managed to remain calm.
Refer to caption Refer to caption Refer to caption
But Tom went on to reveal many unsavory facts about Gatsby. As a result, Gatsby completely flew into a rage. Witnessing this, Daisy also broke down in tears.
Refer to caption Refer to caption Refer to caption
Gatsby still wanted Daisy to leave with him. But Daisy was already in distress and ran out of the crowd. Tom stood there smiling, knowing he had won.
@@
  Andy was wrongfully convicted because he was suspected of murder. The lawyer didn’t believe in Andy’s innocence and sent him to prison. Meanwhile, Red was also undergoing a parole review.
@@
But unfortunately, his parole was denied. After entering prison, Andy and Red formed a connection. Red found Andy interesting and they quickly became friends.
@@
One day, Andy and his friends were working outside. Andy heard the sheriff worrying about economic issues. Ignoring the guard’s objections, Andy seized the chance to offer help to the sheriff.
@@
As a result, Andy won beer for his friends. The warden heard of Andy and had him work for him. Andy agreed.
Refer to caption Refer to caption Refer to caption
Andy took this opportunity to improve the quality of the library in the prison. Inmates’ quality of life became better. But the warden was no longer willing to let Andy leave prison because he knew too much about the warden’s dirty work. Andy told Red to look for something he left under a tree after his release.
Refer to caption Refer to caption Refer to caption
Initially, Red took this as a joke. But Andy eventually succeeded in escaping from prison. Soon after, Red’s parole was also granted.
Refer to caption Refer to caption Refer to caption
He found the oak tree that Andy had mentioned. Beneath the tree, he discovered the letter Andy had left for him. Red was happy because he knew they would meet again.
@@
  Gerald is a math professor, and one day he left a difficult problem on the blackboard. Will is a janitor on campus, and he noticed the problem the professor left behind. Meanwhile, Will is a genius in mathematics, and he quickly became immersed in thought.
@@
At night, Will would go to the bar to hang out with his friends. But after returning home, Will pondered the solution in front of the mirror. The next day, Will left the answer on the blackboard.
@@
The professor was very surprised, but no students claimed they solved it themselves. One day, a student troubled Will’s friend at the bar. But Will managed to refute him with his talent.
@@
This helped him win Skylar’s favor. However, Will is a troubled youth. He was taken to court for fighting. The judge intended to send him to prison.
Refer to caption Refer to caption Refer to caption
The professor was already aware of Will’s situation and proposed that he study mathematics and undergo regular psychological evaluations so he wouldn’t have to go to prison. But many psychologists are insulted by Will. So Gerald had to turn to his old friend Sean for help. At first, Sean also found it difficult to cope with Will.
Refer to caption Refer to caption Refer to caption
But he didn’t give up. His sincerity eventually touched Will. In the end, Will achieved redemption with Sean’s support. After that, Will embarked on a new journey in life.
Figure 13: Additional Zero-Shot results.
Refer to caption Refer to caption Refer to caption
James is checking the spacecraft system. This is his first mission. He is kind of nervous. James launches the spaceship. They are about to land on a new planet.
Refer to caption Refer to caption Refer to caption
Meanwhile, the command center is debating whether they should land on this planet. Steven believes that this planet is worth exploring. However, others disagree with Steven’s idea, as the planet is full of unknowns and could pose dangers.
Refer to caption Refer to caption Refer to caption
But Steven believes that high risks come with high rewards, and he stands by his opinion. They have a heated debate. In the end, Steven convinces everyone and sends the landing instructions to the spaceship.
Refer to caption Refer to caption Refer to caption
James and the astronauts accompanying him receive the instructions and begin preparing for landing. Steven is directing James and the others from the command center as they prepare for landing. The spaceship gradually approaches the planet’s surface.
@@
Figure 15: Additional story comparison results.
@@
To enhance our model’s ability to maintain character consistency, we additionally input the description embedding encoded by LongCLIP (Zhang et al., 2024) and the face embedding extracted by FARL (Zheng et al., 2022) into the trained decoder. Both the text embedding and face embedding are aligned with the decoder’s input using two-layer MLPs and then concatenated with the compressed tokens to serve as the input for cross-attention.
We assume a maximum of 15 sentences for description embedding and 5 character IDs. If the inputs do not meet these quantities, we pad them with zeros. Moreover, for the concatenated embedding, we randomly mask each token with a probability of 0.2. This encourages the model to learn the relationship between facial features, text descriptions, and the compressed token.
@@
The $k$-mixutre GMM models the distribution $p(\mathbf{e}_{t}|\mathbf{H}_{<t},\mathbf{M}_{t})$, where $\mathbf{e}_{t}$ is the continuous real-valued image tokens of sequence length $S$ and dimension $d$. $\mathbf{H}_{<t}$ and $\mathbf{M}_{t}$ refer to the previous history information and the current multi-modal script. The GMM consists of $k$ multivariate Gaussian distribution components that are independent across the channel dimension $d$. To construct the GMM, we replace the logit prediction layer of our autoregressive model initialized using LLaMA with a linear layer with $2kd+k$ dimension for output. Subsequently, the mean $m$ and the variance $v$ of shape $S\times kd$ and mixture weights $w$ of shape $S\times k$ are derived. $m$ consists of a stack of tensors $[m^{(1)},m^{(2)},\dots m^{(k)}]$, where $m_{i}$ is of shape $S\times d$. Similarly, $v=[v^{(1)},v^{(2)},\dots v^{(k)}]$. The mixture weight is normalized along the channel dimension. Given the ground truth image embedding $\mathbf{e}$, we aim to minimize its negative log-likelihood within the GMM:
@@
During training, we additionally introduce the $\ell_{1}$ and $\ell_{2}$ loss between the predicted image tokens and the ground truth tokens to improve the model’s performance. However, we find that these losses are not on the same scale as the negative log-likelihood. The negative log-likelihood is way larger than the $\ell_{1}$ and $\ell_{2}$ loss, which reduces the influence of the $\ell_{1}$ and $\ell_{2}$ loss during the optimization. The difference in scale is because the constant term and covariance matrix of the negative log-likelihood are directly related to the dimension $d$. Concretely, for a variable $\mathbf{x}$ with mean $\mu$, the calculation of the negative log-likelihood is essentially represented as:
@@
where $\frac{d}{2}\log(2\pi)$ is the constant term and $\Sigma$ is the covariance matrix. As $d$ increases, both the constant term and the size of the covariance matrix grow, resulting in a significant increase in the loss value. To bring the different losses to the same scale, we scale the negative log-likelihood loss by dividing it by the dimension $d$. Consequently, the negative log-likelihood loss in Section 3.4 is formulated as: